solve the equation:
5x/4+1/2=x-1/2
thanks and i hope you can help

Answer by ankor@dixie-net.com(22740)

solve the equation:
:
5x/4+1/2=x-1/2
Brackets would be a big help here. Assume you mean:
:
[5x/4] + [1/2] = x - [1/2]
:
Eliminate the denominators, mult equation by 4 and you have:
5x + 2 = 4x - 2
:
5x - 4x = -2 - 2
:
x = -4